OpenStudy (anonymous):

logb(square root x/z^4(y^3) use properties of logarithms to expand the expression?

OpenStudy (anonymous):

\[\log_b{\sqrt{\frac{x}{z^4y^3}}}\] ^Like that?

OpenStudy (anonymous):

apprentices after b, square root only on x, y^3 separated from z^4 to the right

OpenStudy (anonymous):

jabberrwock there?

OpenStudy (anonymous):

Like this? \[\log_b(\frac{\sqrt{x}}{z^4}y^3)\]

OpenStudy (anonymous):

yes definitely :)

OpenStudy (anonymous):

The first thing to do is recognize that \[\sqrt{x}=x^\frac{1}{2}\] Familiar with that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Then the next thing to see is that \[\log(ab)=\log(a)+\log(b)\] So you can rewrite the expression in the following way: \[\log_b(\frac{x^\frac{1}{2}}{z^4}y^3)=\log_b(\frac{x^\frac{1}{2}}{z^4})+\log_b(y^3)\]

OpenStudy (anonymous):

Okay to there?

OpenStudy (anonymous):

it's that it?

OpenStudy (anonymous):

No, not yet

OpenStudy (anonymous):

Then you have to recognize the following: \[\log(\frac{a}{b})=\log(a)-\log(b)\] So you get \[\log_b(\frac{x^\frac{1}{2}}{z^4})+\log_b(y^3)=\log_b(x^\frac{1}{2})-\log_b(z^4)+\log_b(y^3)\]

OpenStudy (anonymous):

Next, we have to deal with the powers, so we use the following property: \[\log(a^b)=b \log(a)\] \[\log_b(x^\frac{1}{2})-\log_b(z^4)+\log_b(y^3)=\frac{1}{2}\log_b(x)-4\log_b(z)+3\log_b(y)\]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so that's it? I will need to practice a lot to get that through my head. lot of steps but it's worth it.

OpenStudy (anonymous):

Yep, as long as you use one property at a time, it should work out.

OpenStudy (anonymous):

another question

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

Condense the expression to the logarithm of a single quantity. 1/3[lnx+3ln(x-2)]-ln(x+1)

OpenStudy (anonymous):

You're going the other way with this one. \[\frac{1}{3}(\ln(x)+3\ln(x-2))-\ln(x+1)\] Like that?

OpenStudy (anonymous):

first x don't need apprentices or maybe it don't matter

OpenStudy (anonymous):

Shouldn't matter.

OpenStudy (anonymous):

Alright, just like always, we're going to start with the innermost set of parentheses.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Let's start with the part that says 3ln(x-2) Use the property \[b \ln(a)=\ln(a^b)\] We get \[3\ln(x-2)=\ln((x-2)^3)\]

OpenStudy (anonymous):

Next, we look at this part: \[\ln(x)+\ln((x-2)^3)\] We use the property \[\ln(a)+\ln(b)=\ln(ab)\] And we get \[\ln(x)+\ln((x-2)^3)=\ln(x(x-2)^3)\]

OpenStudy (anonymous):

So, basically, those two logarithms became one larger logarithm.

OpenStudy (anonymous):

Here is where we're at now: \[\frac{1}{3}(\ln(x(x-2)^3)-\ln(x+1)\]

OpenStudy (anonymous):

Now we use the property \[b \ln(a)=\ln(a^b)\] We get \[\ln((x(x-2)^3)^\frac{1}{3})-\ln(x+1)\]

OpenStudy (anonymous):

big steps

OpenStudy (anonymous):

Then, finally, we use the property\[\ln(a)-\ln(b)=\ln(\frac{a}{b})\] \[\ln(\frac{(x(x-2)^3)^\frac{1}{3}}{x+1})\]

OpenStudy (anonymous):

Yeah, need me to go through any of them again?

OpenStudy (anonymous):

no I have a next problem give me a minute

OpenStudy (anonymous):

approximate the following logarithms. a) logb9 b)logb4/9

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

The given are logb2=0.3562, logb3=0.5646, log5=0.8271